OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-8, -11) and is perpendicular to -2x – 7y = -14 Please show all of your work.

OpenStudy (jlastino):

First find the slope of -2x-7y=-14 (change it into the form y=mx+b) -7y=2x+14 y=-2/7x-2 So the slope of this line is -2/7 BUT this line is perpendicular to the line being ask so we take its negative reciprocal to find the slope of the line we are looking form 7/2 Now we will use 7/2 and (-8,-11) to find the line using the point slope equation y-y1= m (x-x1) where y1=-11 x1=-8 m=7/2 So y+11=7/2(x+8) y+11= 7/2x +28 y=(7/2)x +28-11 y=(7/2)x +17 <---slope intercept form

OpenStudy (anonymous):

Thank you